OpenStudy (anonymous):

\[f(x)=3(x-1)^{2}-2\] x-intercept answer has to be a decimal??? help!

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

and there are two of them

OpenStudy (anonymous):

\[x = 1 +- \sqrt{ \frac{2}{3}}\]

OpenStudy (anonymous):

what would y-intercept be for this???

OpenStudy (anonymous):

1.816 and 0.1835 are two x intercepts

OpenStudy (anonymous):

set x=0, for y intercept, y= 3 (-1)^2 -2 = 3 -2 =1

OpenStudy (anonymous):

axis symmetry?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

thats what i thought but the computer is telling me that's wrong